If the marginal cost of making a photocopy is 3 cents and the elasticity of demand is –2, the profit-maximizing price is:

a.

5 cents

b.

8 cents

c.

6 cents

d.

3 cents

e.

7 cents
Answer #1

Ans) Profit maximising price, P = MC [(e)/(e+1)] ; where e is elasticity

P = 3 [(-2)/(-2+1)] = 3× (-2/-1) = 3×2 = 6cents.

Option c.